OpenStudy (anonymous):

what is x+11=3x+3 x=?

OpenStudy (anonymous):

First subtract 3x from both sides which give -2x + 11 = 3 Next subtract 11 from both sides which gives -2x = -8 Finally divide both sides by -2 which gives x = 4 The trick to solving algebra problems is combining like terms

OpenStudy (anonymous):

x+11=3x+3 you would subtract x from both sides first 11= 2x+3 then you want to get the numbers on one side of the equation so you subtract 3 from both sides 8= 2x divide 2 from both sides 4=x

OpenStudy (anonymous):

x+11has to equal the same as 3x+3

OpenStudy (anonymous):

what is x+3+3x=x+15 x=?

OpenStudy (anonymous):

well, what value do you think it is? if x = 3, the left is bigger than the right if x = 5, the right is bigger than the left there is SOME number between there that makes them equal...